OpenStudy (anonymous):

type the formula for the number of combinations of n things taken r at a time. c(n,r)

OpenStudy (anonymous):

\[\Large \left(\begin{matrix}n \\ r\end{matrix}\right)=\frac{n!}{r!(n-r)!}\]

OpenStudy (larseighner):

\( \huge {n \choose r }= {{n!} \over {r!(n−r)!}} \)